Circular No. 6729

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970828 A. V. Filippenko, D. Stern, and R. R. Treffers, University of Califor-

nia at Berkeley; C. Y. Peng, University of Arizona; H. E. Bond, Space Telescope Science Institute; and M. W. Richmond, Rochester Institute of Technology, report that a total of ten R-band images (size 5′ × 5′) centered on the nominal position of GRB 970828 (cf. IAUC 6726) were obtained on Aug. 29.3 UT at Lick Observatory, using the 1-m Nickel Telescope and the 0.8-m Katzman Automatic Imaging Telescope. No new objects brighter than the limiting magnitude of the first Palomar Observatory Sky Survey (digitized or photographic) are visible in the images, which span the 90percent-confidence ellipse given on IAUC 6728. The field will be monitored for variable objects over the next week.

T. Murakami, Y. Ueda, and M. Ozaki, Institute of Space and Astronautical Science; A. Yoshida and N. Kawai, Institute of Physical and Chemical Research; M. Yamauchi, Miyazaki University; and F. E. Marshall, R. H. D. Corbet, and T. Takeshima, Goddard Space Flight Center, report on behalf of the ASCA team and the RXTE Science Operations Team: “ASCA is observing the error box of GRB 970828 following the detection by RXTE’s ASM and PCA (IAUC 6726, 6727, 6728) from Aug. 29.91 to 30.85 UT. A preliminary analysis of about half of the total data set reveals an x-ray source that is consistent with the combined error box of RXTE/ASM and IPN (IAUC 6728). The preliminary position of the x-ray source is α = 18^{h}08^{m}30^{s}, δ = +59^{o}19′20′′ (equinox 2000.0; error radius 1′.5). Analysis of the full data set after the attitude determination will improve the position and reduce the error. The results will be reported later.”

E. Brocato, Osservatorio Astronomico Collurania (OAC), Teramo; M. Stiavelli, Space Telescope Science Institute (STSI), European Space Agency, and Scuola Normale Superiore, Pisa; and H. E. Bond and K. Sahu, STSI, report that a total of four 30-min R and I band CCD images (size 4′.5 × 4′.5), spaced to cover the 90-percent-confidence ellipse at the position of GRB 970828 (IAUC 6728), were obtained at the 0.7-m OAC telescope on Aug. 29.84–29.99 UT. No new objects brighter than the limiting magnitudes of the first and second Palomar Observatory Sky Surveys are visible in the images.

1997 August 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green